This logbook is in reverse entry format. The first entry is at the bottom.
12/18/2015 - Well after 6 months the observatory has finally been picked up by
the Kennesaw Mountain High School science department. It will now be in storage
at the school until it is ready for installation. I have offered to help them
get it up and running.
7/26/2015 - The observatory has been disassembled and crated due to our move to
Georgia. I am in the process of trying to donate it to a local Georgia high school. The new place has way to many trees.
10/27/2014 – Replaced the ProStar Crated Observatory
batteries. The new controller is an upgrade over the old one as it has an LCD display for Batt volts,
Solar amps and batt amps. The two 115 ah batteries were replaced with one 122 ah battery as the
store only had one.
- Installed 800 watt inverter with remote control.
6/26/2014 - Reconfigured telescopes.The configuration is now:
2.6mm Tamron Wide Angle Lens - Imager = GStar-Ex B&W Video.
40mm Coronado Solar Spoce - Visual Onl
50 mm DIY Reflector - Imager = SBIG ST-2000XM with 10 slot filter wheel
66mm Williams Optics Refractor - Visual Only
102mm Stellarvue APO Refractor - Imager = SBIG ST-402ME with LRGB cfw.
- Both focuser ports on the Paramount ME stopped working. After reseting the Bisque
TCS file parameters back to the default values the focuser ports works again.
- Assembled a DIY 50mm refractor for use as a wide field imager. The scope will be
attached to the SBIG ST-2000XM. The scope has a motorized focuser and was constructed
from a Meade 50mm finder scope.
4/26/2013 – Washed dome. Tightened acorn wheel nuts.
4/24/2013 – Installed the 10" Meade SCT. The telescope configuration is now:
2.6mm Tamron Wide Angle Lens - Imager = GStar-Ex B&W Video.
40mm Coronado Solar Spoce - Visual Only
50mm Stellarvue APO Refractor - Imager = Mallincam Hyper Color Video
66mm Williams Optics Refractor - - Visual Only
102mm Stellarvue APO Refractor - Imager = SBIG ST-402ME with LRGB cfw
254mm Meade SCT Reflector - Imager = SBIG ST-2000XM with 10 slot filter wheel
3/05/2013 – Repaired the Boltwood Cloud Sensor. The thermal grease used to both
seal out the elements and provide electrical isolation of the IR sky temperature
sensor and the rain sensor circuit needed to be replaced. The sensor would always
report the sky temperature as being between -10F an +4F. That would cause the
software to always display "Very Cloudy" and constantly trigger the shutter close
routine for the dome. The sensor will be placed back online on the website's
weather page in a few days, after it is fully calibrated.
2/21/2013 – Replaced the Davis Vantage Pro 1 data logger. Weather Station CW2775
is again uploading data to the website and Citizen Weather Observer Program.
9/29/2011 – The observatory has been completely upgraded to TheSkyX and all the
current available software upgraded such as TPoint, CCDSoft etc.
- Moved the sensor for dome rotation from the motor shaft to the rotation gear.
A black and white target was painted on the gear. This setup is much more
reliable and corrects an issue with the dome rotation and slit alignment to the
- Reconfigured the telescopes to the 10” SCT with a piggybacked 66mm Refractor.
- Observatory control program modified version 10.1 in order to support TheSkyX.
- Petcool configured for winter (heater on A/C off).
- Weather station remains off-line due to lightning strike.
- Experimented with POTH telescope hub. Dome alignment accuracy not so good.
- Mounted 1200 watt inverter on south wall and wired it into battery box.
Reorganized power cords so only one AC plug needs to be moved in order operate
the observatory from battery pack during a mains power outage.
1/9/2010 - Continue to work issues with upgrading the observatory to TheSkyX.
- Ordered camera extension cable from SBIG.
12/29/2010 - Began configuring the mount for the SV102A, SV110ED and the
ZenithStar 66 in a Side-by-Side arrangement. The SV102A has the ST-2000XM, the
SV110ED has the ST-402ME. ZenithStar 66 has the Mallicam attached. The setup
weights 60lbs without the SBIG cameras attached.
12/10/2010 - Installed TheSkyX on the observatory computer for testing. TheSky6
remains the primary control software for the Paramount ME.
11/21/2010 - Reconfigured Petcool for winter heating.
weather station power control off-line due to unknown issue
6/14/2010 - Installed the 12" LX200R and the ETX90 OTA.
- Received the ST402-ME back from repair. Not tested yet.
3/14/2010 - Removed the 10" LX200GPS OTA and cradle.
- Installed the ADM side by side plate with cradles for the SV102 and 110.
The pointing adjustment plate was attached to the 110 cradle.
- Fix the rain gage, a twig was caught in the spoon. The debris mesh is missing.
- Adjusted the All SkyCam script for daylight savings time.
12/24/09 - The Sky Quality Meter has been repaired. As suspected the power plug
had come loose.
12/24/09 - The Sky Quality Meter has not worked since the snow storm on Friday
night. A quick inspection this morning shows some slack in cable. I believe the
power cable has become disconnected internally. I will open the unit later for
12/21/09 - Called SBIG about my ST-402ME. Again my call to the repair department
rang over to voice mail. This is the 2nd VM I have left along with two emails.
There have been no response to either my VM or emails. I am not happy with SBIG
at this point.
10/12/09 - The observatory has been indefinitely idled.
- Petcool is in heat mode and being operated via the humidity controlled power
- Computer, monitor, mount and all cameras are disconnected and/or powered
- Shutters are locked out via the internal switch.
- Scopes are capped and covered.
- Pier lowered
8/22/09 - The SkyCam is up and running. The problem turned out to be a lose
8/22/09 - The SkyCam is not working. The problem has not been determined.
- Cameras and telescopes reconfigured:
254mm SCT w/SBIG ST-2000XM
66mm Refractor w/Mallincam
80mm Refractor w/SPC900NC/27
- Base of wall and deck interface sealed with caulk.
8/8/09 - The ST-402ME has been shipped back to SBIG for repair.
8/3/09 - Installed the Optec adjustable Az/Alt dovetail plate.
The ST-402ME suffered a electronics failure (bright flash & smoke). it has been
replaced with the 2nd ST-2000XM with CFW5.
6/20/09 - ASCOM Dome Parameters for Side-by-Side configuration
SV110 & SV102
latitude : 40.16
Pivot Point East (inches): 6
Pivot Point North (inches): -6
Pivot Point Up (inches) -10
Pivot Point Optical (inches): 18
Tic Per Revolution : 228
Home: 0 degrees
Park: 325 degrees
6/16/09 - Removed spectrograph and the 12" SCT. The configuration is now the
110mm and the 102mm refractor side by side.
5/20/09 - Repaired the wireless shutter operation, a loose screw had allowed
the a wire to become disconnected.
5/19/09 - Attached the SV110ED atop the LX200R. The custom ADM side-by-side
plate was also attached. The WO-66 with the ST2000-XM is on one side and the
GStar-EX with the 2.6mm lens on the other.
5/17/09 - Installed the LHIRES-III spectrograph in the 12" LX200R OTA @ F10.
The SBIG ST-402ME is the attached to the imaging port and the Meade DSI-Pro is
connected the guider port. CCDSoft has be setup for the cameras.
5/16/09 - The dome rotation issue turned out not to be a motor or controller
board issue. The problem was a mis-aligned optical sensor. Once the sensor was
realigned the issue was resolved.
- Moved the 2 port PCI serial card to the bottom PCI slot. The dome
rotation & mount control are working again. Was not able to load the drivers
for the MosChip 4 port card on the observatory computer. This card installed
without issue on the Weather Station PC. I suspect the issue may be related
to Windows SP3.
- Sealed the based of the observatory with polyurethane in an attempt to keep
the water out.
5/14/09 - Received the 300 and 1200 LHIRES-III grating modules.
5/1/09 - Received and installed the replacement LHIRES-III grating module.
4/29/09 - Installed a 2 TB Cavalry CAND Raid 1 network storage drive.
4/28/09 - The complete 2400mm grating module was returned to Adirondack for a
4/27/09 - The 2400mm grating mirror detached itself from its module plate.
- Purchased a Shelyak LHIRES-III spectrograph with 2400 & 150mm gratings. When the
LHIRES-III is used in stand alone solar mode, it will be mounted on the LXD-650.
- All telescopes have been removed from the observatory.
- Dome motor has been swapped out. The dome now gives the following error
during rotation: "Communications Error"
- Begun the swap out of the dome rotation motor.
- Replenished water level in the 220ah battery pack. The batteries required
almost 1/2 gallon of water to top them off.
- Replaced the Dec motor and control board in the LX200GPS. While testing the
scope today it reported a motor stalled error.
Pre and post repair pictures of the Dec assembly can be found here:
- Swapped MaxDome-II dome rotation board serial #320 with board #1087.
- Began cabling up the new scope configuration. The setup is very closed to
being balanced. None of the scope have be focused yet.
The image below is from the GStar-EX mount between the scopes. It will be aligned
when the scopes are focused.
- Removed the Meade 12" Meade SCT from the mount.
- Installed the Stellarvue SV110ED-A and SV102A side by side. The Williams
Optics Zenithstar 80mm has been piggy-backed on the SV102A and the Coronado
ScopeMax 40 hydrogen Alpha solar scope is piggy-backed on the SV110ED-A. The
- Installed the remaining 7 dome wheels. The 8 additional wheels put the spacing
between them at 17.5 inch. That is half of the original 35 inches from the
manufacturer. The picture below shows one of the new wheels (center).
The hole was cut to allow wheels to be installed or removed without removing the
dome. (The hole is not visible when the wheel cover is in place.)
- Installed one of the dome wheels. Originally the wheels were spaced 35 inches
apart. The new wheel was place in the center of the area were the lower part of
the dome was dragging on the top of the walls. A noticeable improvement in the
smoothness of the rotation was observed.
- The dome wheels and motor arrived a few day ago, but it been a bit to chilly
for me to do the installation.
- Update the firmware in the SQM-LE to 3-3-12. This allows the calibration
values to be changed. The jSQM-LE.jar program was also updated.
- Installed the SQM Reader Pro software on the weather station PC. The new
software provides graphing and FTP ability.
- Yesterday was a good and bad day. I decided to drill the vent holes in the
PVC that holds the SQM. This will help equalize the internal and outside
ambient temperatures. Without the holes the internal temps are 20+ degrees
higher than outside. I decided to leave the SQM circuit board in place during
the drilling. Bad move. On the last hole the drill bit caught and damaged some
traces on the circuit board. The good news is I was able to repair the traces
with some bus wire and a soldering iron. The other good news is the vent holes
work. The the temps inside are only about 4 degrees above ambient. I
will have to keep a close eye on the temps in the cold weather. If the temp
drops too low inside, condensation could build up on the inside cover. To correct
that the design allows for the top and middle pieces of PVC to be rotated to
close the holes thereby raising the internal temps.
- It appears I may have found the issue with the dome rotation. The dome is
dragging on the top of the dome walls in some places. I am not really sure
why this is happening. It is almost as if the dome is distorted. In order to
fix the problem I have ordered a set of 8 wheels from the dome manufacturer.
The wheels will be added as needed to the low spots on the dome. The existing
wheels will also be removed, cleaned and reinstalled. A 3/4" hole was drilled
on the inside of the dome track near the man door. This hole will allow access
to the screw head that holds the dome wheel on the bearing. Without the hole
the dome would need to be lifted off the walls to access the screw. The hole
will be cover when that section of the protective wheel cover is put back in
place. That section of the cover will be marked so the location of the hole can
easily be located if access is needed in the future.
- The outer surface of the top of the dome walls and the inside of the rubber
weather strip were recoated with silicone lubricant.
- Ordered a spare dome rotation drive motor and spare drive gear.
- The SQM-LE has been permanently mounted on the bottom (east side) of the 120
watt solar panel. The initial mounting position was adjusted lower as it was
casting a shadow on the panel. The meter was also recalibrated in hopes of
producing a more reasonable reading of the local sky darkness. Prior readings
seemed to show the sky was darker than it really was.
- The SQM meter is now being tested outside. It is not yet permanently mounted.
As soon as the weather warms a hole will be dug for the PVC pole it will sit
on. The enclosure info is here. http://jatobservatory.org/SQM/SQM_kit.html
- The dome rotation issue has not been fixed.
- The observatory's dome stopped rotating on 1/16/09. The dome stopped while
it was facing SSE. The 5 amp fuse inside the dome control panel was blown. The
fuse was replace and the dome is able to rotate again, but the rotation is not
smooth. It appears the dome is binding or dragging so the observatory has been
- The Unihedron Sky Quality Meter (SQM-LE) arrived today. The meter uses a
Ethernet port to make sky darkness readings. The folks at Knightware (the makers
of Deepsky4) provide a free utility to read the meter. Their SQMReader program
displays the SQM output in magnitude per square arcsecond (MPSAS) and Naked Eye
Limiting Magnitude (NELM).
Greg Kettell from Cloudynights was kind enough to
write some code for me in php that reads the output file from the Knightware
software and display the data in a readable and useful format.
You can see the output here: http://jatobservatory.org/SQM.php Currently the
data is only test data and doesn't represent actual sky darkness values. Also
since the SQM-LE is not weather proof it must be placed in an custom enclosure,
which I have yet to finish the design.
I am also waiting for the Extech light meter to arrive since I managed to over
write the calibration file in the SQM when I connected it the PC (always read
the manual first).
- Received 2nd Petcool HAC unit. This will be a spare
- Received Orion V-Block, Sulfur-II and Hydrogen-Beta 1.25" filters
- Removed Schuller VBRI filters for the CFW10
- Installed the new filters in the following positions:
Hydrogen-Beta filer in filter wheel slot #7 of ST-2000XM #1
Sulfur-II filer in filter wheel slot #8 of ST-2000XM #1
UltraBlock LP filer in filter wheel slot #9 of ST-2000XM #1
SkyGlow LP filer in filter wheel slot #10 of ST-2000XM #1
- Cleaned all filters and the 6.3 Focal Reducer
- Modified OCP8 replaced CCDCommander button with CCD Navigator.
- The plate solve issue experienced a few days ago was a result of the
astrometry catalogs being turned off within TheSky. The catalogs can be turned
back on via the tool bar by selecting "View">"Stellar Options">
- Fixed some issues with MS Agent that were preventing ACP from running after
the system restore.
- Modified OCP to ver 8.3.4 to include TPoint Mapper and the ASCOM Dome
controller. The code will terminate ACP when TPoint Mapper starts. It will
also terminate ASCOM Dome when ACP is started.
- Created new batch files TPointMapper.bat, ASCOMDome.bat and
Kill_TPointMapper.bat to handle the starting and stopping for the new code.
- Configured TPointMapper with a 208 point model of the eastern sky. Run was
aborted due to clouds.
- Created a new system backup of the observatory computer. This one includes
all fixes and mods listed above.
- Created a system backup of the observatory computer.
- Created a 208 point TPoint model using AAGware's TPoint Mapper software. I
started using the TPoint Mapper software because of the dome lag issues my
system suffers from when not using Automadome as the dome controller. I also
have "out of memory errors" when using Automapper 2 and the mapping runs never
complete. The mount will need to have its Polar Alignment readjust based on the
data returned by the TPoint run:
************* AZIMUTH ************
MA: +46 seconds (+0.8 minutes).
Rotate axis West (counterclockwise).
For latitude 40.17°, the azimuth adjustment is 1.0 minutes.
Paramount ME: Loosen the West knob, and tighten the East knob 0.5 knob tics
************* ALTITUDE ************
ME: +709 seconds (+11.8 minutes).
The polar axis should be raised +709 seconds (+11.8 minutes).
Paramount ME: Raise the polar axis 5.9 knob tics
- had an issue where the CCDSoft would no longer plate solve. I was getting
error messages ranging from things like"to few stars" in image to "error unknown".
I know the images where fine because I have what I call a plate solve calibration
image that I know works. After explore all options that including loading a
previous day's restore point so I reloaded the system. That fixed the plate solve
issues, but it broke ACP.
- Reinstalled ACP and applied latest hotfix.
- Current ASCOM dome parameters:
Latitude : 40.16
Pivot Point East (inches): 2
Pivot Point North (inches): 1
Pivot Point Up (inches) -25
Pivot Point Optical (inches): 20.276
Tic Per Revolution : 227
Home: 0 degrees
Park: 336 degrees
- Bob Denny from DC3 Dreams the author of the ACP Observatory Control software
was kind enough to create a script for ACP that using the Boltwood Cloud sensor
data file. Previously ACP could only make use of the Boltwood senser is it was
connected to observatory computer via a comm port. Mine is not so this feature
is very useful to me. Thanks Bob.
- Installed the 2.5 amp power adapter for the Mallincam and GStar CCD video
cameras. All 60 cycle noise in gone and the cameras start up faster.
- The observatory now supports web based scheduling via ACP's "Scheduler"
software. The simple logs in to the web scheduler and enters the target,
filters, Ra, Dec, expose time and binning. The request is submitted to the
observatory and a run automatically scheduled.
- Received the Matrix CCD Video camera from the CN Swap & Shop purchase. The
camera works but has a lot of hot pixels. It also draws more power the either
the Mallincam of GStar camera. A new power supply will have to be installed in
the observatory before the Matrix camera can be used as a replacement. The
Matrix camera does not support remote control of the settings.
- The 60 cycle ground issue turns out to be the 500ma power supply used to power
both the Mallincam and GStar CCD video cameras. The powers supply appear to
not to putting out the power needed to drive both camera at the same time.
A replacement power which supplies 2.5 amps will be installed. Until then the
Mallincam with remain unplugged.
- Most of the 60 cycle noise problem was due to the Peltier cooler on the
Mallincam. When the cooler was turned off most of the noise disappeared.
- Removed the Moonlite focuser and replaced it with the Meade Micro-Focuser.
The Moonlite unit was slipping, causing focusing isses for the SCT.
- Received and installed the mini fuse for AUX port 1 in the Paramount ME.
The fuse controls the dew heater, It turns out the fus was not blown but the
AUX-1 connector on the Vesa-Plate side was cross threaded and resulted in a
an open circuit condition to the dew controller. But now I at least have
space fuses. They are Little Fuses part number: 0452003.MRL. They were ordered
from Mouser under the Mouser part number 576-0452003.MRL. The fuses are 3 amp.
- The observatory has developed a 60 cycle noise issue on both the MalinCam and
the GStar-Ex video cameras. It suspect there is an grounding issue with the
power supply (both cameras are fed by the same power supply).
- Added 1 gig of memory to the obs PC for a total of 2GB.
- Restored the system for Backup.
- Modified the OCP program to include: SBScriper, MPO Connects, MPO AVG, IRIS
CCDSoft, K3CCDTools, CCDAutopilot4 and CCD Commander. Program now turns on
the heater since it is getting cooler. K3CCDTools is noe the default video
monitor when the sysem starts up.
- Removed and cleaned the ST-2000MX CCD camera again. Dust was present on either
the CCD chip, cover glass or filter(s).
- Since the main camera was removed I realignment was needed. The home index
offset was reset and a new T-Point model was created. The model suggested the
following adjustments to the polar alignment:
************* AZIMUTH ************
MA: +1451 seconds (+24.2 minutes).
Rotate axis West (counterclockwise).
For latitude 40.17°, the azimuth adjustment is 31.6 minutes.
Paramount ME: Loosen the West knob, and tighten the East knob 1.1 full turns
************* ALTITUDE ************
ME: -596 seconds (-9.9 minutes).
The polar axis should be lowered -596 seconds (-9.9 minutes).
Paramount ME: Lower the polar axis 5.0 knob tics
- Cleaned the #1 SBIG ST-2000MX CCD window and filter wheel.
- Installed software as of 8/24/08
- The observatory is back on-line and fully operational. The problem turned out
to be a faulty USB port that was dragging down the 12 volt power supply.
- USB routing was redone so the 2 USB hubs where not in a chain. They are
now connect to root ports.
- Current Com port configuration:
Com 1 = Serial Switch Box - connected to onboard serial port
Com 2 = Power control outlets - connected to USB to serial adapter (USB hub)
Com 3 = Paramount ME - connected to left port on PCI board
Com 4 = Dome rotation - Connected to right port on PCI baord
Com 5 = reserved for camera rotator - USB to serial adapter (USB hub)
Com 6 = Undefined - connected to USB to serial adapter (USB hub)
- Installed MaxIm DL 5.00
- Toped off the water in the battery pack. Both were low.
- The observatory computer suffered a boot failure during the driver loading.
The failure occurs after the Mup.sys driver is loaded. The last time his occurred
the power supply had failed (see 2/16/08 entry). The problem appears to be with
either the USB to serial converters or the 2 port serial PCI cards. The computer
was removed from the observatory and the PCI serial card was removed. The BIOS
was reset and the computer was able to be booted to the Windows OS. The current
problem is the serial ports (both PCI and USB) are either defective or not
configured correctly as communication with the scope is not possible on com
port 5. Each com port will have to be traced and tested.
- Installed the Shorty metal dew. This dew was originally designed for the 12"
RCX400. The dew shield also has a triangular Hartman mask.
- Aligned the 3 telescopes SCT, 50mm and 66mm so all are pointing to the same
- Created a T-Point model.
- Modified the Center Image script used in CCDSoft to display crosshair. The
script will now turn on the crosshairs, center the brightest star (or object)
then turn off the crosshairs.
- Installed the Willams Optic 66mm refracto attached the SBIG ST-402ME camera
- Installed the Stellarvue 50mm F50e finder and Mallicam video camera.
- Removed the Stellarvue SV102APO refractor
- Removed the Pyxis camera rotator
- Removed the Meade micro focuser
- After rewriting and testing the Observatory Control Program, version 8.3.2
is ready for full time use. The new version cleans up the code but has no new
Pyxis camera rotator
Meade micro focuser
- Removed the telscope "array"
- Installed the 12" SCT, moonlite focuser and Williams Optics 66mm refractor.
- Attached the SBIG St-402ME to the WO66mm.
- Attached the Mallincam to the SCT.
- Attached the Stellarvue 50MM finder to the SCT
- Attached the Telrad to the SCT
- Installed dew heaters on the SCT, 66mm and and 50mm finder.
- Balanced the setup
- Received the 12" LX200R OTA, Moonlite focuser, Rigel Systems nFocus controller
dew heater, Telrad and SCT ventilator.
- Received the ST-402ME back from SBIG. The camera was sent in to repair a
no connection issue. SBIG replaced a board at No Charge.
- Installed the 10" OTA back on the forks. The optics still need to be
collimated. Handbox version is 4.1i.
- Added a blog to the website.
- Added a new inlet to the heating and A/C for the observatory. The inlet is
just below the rotational dome section. The cool forced air will enter from
there and should allow better cooling. In the winter the warm air will be
drawn from there in the hope of providing better circulation.
- Sent the ST402-ME back to SBIG for repair (unable to connect to the camera).
- The telescope configuration was recently changed. The 10" SCT is no longer
in use. The new configuration consists of 6 refractors ranging in size from
40mm to 102mm mounted on a common fixture. The setup has been nicknamed "the
" and it can either be used for visual observing or imaging. The
GStar-Ex camera that is used for monitoring the dome position is part of the
Array. All of the scopes except for the larger 102mm mounted in the middle
ride on dovetail adapter plates and are adjustable. This allows the Array to
be collimated so all scopes point at the same target as the 102mm.
- The DC to DC power supply for the ST-2000XM camera arrived today. The camera
now operates fine.
- Requested that Bisque provide the new cover for the upgraded MKS-4000 board.
- ST-402ME will be package up and returned the SBIG for repair. Can't connect
to the camera.
- Installed the MKS-4000 controller and interface panel board. Slew
temperature compensation has been set to 50F (10C). New values for
the MKS4000 are below.
Bisque Telescope Control System (TCS)
Copyright © 2002-2006 Software Bisque, Inc.
Mount Id.: Paramount ME Running at: 3333.333
***************** Axis 0 (RA) Control version: 1.1.45 *****************
Index angle: 361
Base rate: -43067265
Maximum speed: 600000000
Non-sidereal tracking rate: 0
Minimum position limit: -4960000
Maximum position limit: 1120000
Sensor HA (from sync): 1.9859
Tics per revolution: 11520000
PEC ratio: 10
Maximum position error: 1000
Unit Id.: 100
EMF constant: 72
Home velocity high: 660000000
Home velocity medium: 110000000
Home velocity low: 3900000
Home direction, sense: 1100
Home mode, required, Joystick, In-out-in: 2111
Home index offset: 0
PEC cutoff speed: 0
Maximum voltage: 15
Maximum gain: 150
***************** Axis 1 (Dec) Control version: 1.1.45 *****************
Index angle: 1369
Base rate: 0
Maximum speed: 600000000
Non-sidereal tracking rate: 0
Minimum position limit: -5460000
Maximum position limit: 1700000
Sensor Dec (from sync): 0.0886
Tics per revolution: -7500000
PEC ratio: 10
Maximum position error: 1000
Unit Id.: 100
EMF constant: 72
Home velocity high: 660000000
Home velocity medium: 110000000
Home velocity low: 3900000
Home direction, sense: 1100
Home mode, required, Joystick, In-out-in: 2111
Home index offset: 0
PEC cutoff speed: 0
Maximum voltage: 15
Maximum gain: 150
- A new power supply for the ST-2000XM was ordered.
- Re-installed the dehumidifier and configured the Heating and A/C unit for
- Re-installed Quicktime 7.1.3 along with new SiS651 video drivers and
Starry Nights Pro now works.
- The LX200GPS OTA needs to be collimated.
- Modified the OCP8 code to turn on/off the dehumidifier. Since the
dehumidifier replaces the Heat & A/C unit the code now also turns off the
A/C when the shutters are open. The interior lights have now been coded to
work on the Aux outlet. The lighting outlet was allocated back to the
- The observatory is off line for some spring maintenance and upgrades. The
control board for the Paramount ME will be upgraded from the old 3000 series
to the newer MKS-4000 series. An attempt will be made to seal the gap where
the upper and lower shutters overlap. New sealant will be applied to dome walls
and decking. Cameras will also be moved between scopes as I continue to have
problems with the SBIG cameras. The ST-2000XM is on it way back from SBIG. It
was returned to them again after repair for a black image. They can't seem to
find a problem. The ST-401ME was sent in for repair for not being able to
connect to the camera. The problem has returned. The camera will have to be sent
- Installed one of the 4 port PCI serial port cards in the Weather Station PC.
The system is functional and data transfer between the weather module in the
observatory and the WeatherLink 5.6 software is much faster. The rain and cloud
detector has also been reconnected to the weather station PC. Both USB to
serial adapters have been uninstalled. There is one open com port (Com5).
- Ordered two 4 port PCI serial port cards. One for the observatory and one for
the weather station PC.
- Repaired broken links on weather page.
- Installed a new 500 watt power supply. The observatory is up and running.
- The main observatory computer has failed. The power supply is burned out.
A replacement supply was installed boot the system would not power up. The
observatory will be offline until the computer is either repaired or replaced.
- The MKS4000 upgrade controlled for the Paramount ME will be installed during
the down time.
- The short belt in the RA axis will also be replaced.
- The ST-2000XM was shipped back to SBIG today.
- The scope's park position had the be reset.
- Received the ST-2000XM back for SBIG. The camera still has the original
problem of a black screen after image download. The ADU count is always 0.
Needless to say I am not happy with SBIG's repair service.
- Ordered the MKS 4000 upgrade for the Paramount Me. SB threw in two new belts
for the Dec drive.
- Replaced the 2 deepcycle 100 amp hour marine batteries, with two 115 amp
hour units. The old batteries had begun to fail.
- Strung a temporary composite video cable between MallinCam and the Dazzle
video capture box. The original cable was damaged and needs to be replaced.
- Aligned the MallinCam, finder scopes and the SV102APO to each other.
- Solved a long standing issue with the dome rotation unexpectedly stopping
during operation. Two switches on the wall mounted rotation interface box had
been left in the wrong position. During remote control operations the wireless
switch box enables the "Dome Enable" and "Dome Power" functions via relays on
the circuit board, but the manual switches must be in the up position.
- The 5 port USB hub on the pier is intermittently cycling off and on during
the dome rotation. I suspect there is an issue with power supply that is being
used to power the dome.
- The Adirondack Flat Fielder light box was remounted on the ceiling just past
the rear of Zenith. The scope and dome where tested to be sure they don't
interfere with each other.
Water was found on the floor of the observatory near the entrance door. It
appears the caulking at the base of the dome's walls where damaged when the
decking was leveled a few months ago. New caulking will be applied to the
full diameter of the wall and decking interface.
- The repaired ST-402ME has returned from SBIG. The camera now has enchanced
cooling. Previously the camera was only able to be cooled to above -3C. Now it
is able to be cooled to about -15C. The ST2K is not expected back until
sometime after the New Year.
- Removed the dehumidifier from the observatory. The drain hose remains for
- Received the Williams Optic 66mm SD 388mm focal length APO refractor.
- Wired up the outside batteries.
- Connected the Velleman voltage/data logger as follows:
Channel 1 - Solar Panel Voltage
Channel 2 - Load Voltage (battery output voltage)
Channel 3 - 12 volt DC main power supply
Channel 4 - Pier Power (DC voltage)
- The #2 analog voltmeter was reconfigured to display the Main DC power supply
voltage instead of battery voltage.
- Shipped the ST-402ME back to SBIG for repair. No info on when either it of the
ST-2000XM will be returned.
- Install the DSI-Color and DSI Pro CCD camera in place of the ST-2000XM and 402.
- Moved the batteries outside the observatory. They are installed in a plastic
Stanley Tool box. They have not been wired to the solar panel and charge
regulator yet. The batteries have been replaced with a 15 amp DC power supply.
This power supply drives the dome rotation only. The 10 amp DC supply is used
to power the pier. During the rewiring some issues have cropped up:
The SBIG 402ME has no power.
The 4.5v power supply smoked due to a voltage reversal during the rewire.
The USB Vellman data logger is not connected.
The Red & White florescent are not working.
-The Observatory Control GUI (ver 8.2.2) has been upgraded and tested.
-Drawers (2 on each side) have been installed to support the computer self.
-The batteries where filled with water.
- The ST-2000XM was shipped to SBIG. It should arrive there on 11/26.
- The Observatory Control Panel GUI is being modified to handle the Heating &
Air Conditioning unit. The new GUI is shown below. A additional remote control
power strip was installed to support automation of the HAC unit. The GUI also
now supports the Sling Player and has the option to reboot the observatory PC.
- Comet 17P/Holmes continues to be an amusing object in the Northeastern sky.
This evening I used my iPhone running the latest version of VNC the control
the observatory and capture an image of Holmes. The first image below shows
the iPhone connected to the observatory computer with is running CCDSoft.
Hers is the actual picture. The image was taken with a Stellervue 104 mm apo refractor
and a SBIG ST-402ME. Until the ST-2000 I returned to SBIG for repair comes back the
402 ME is my primary imager.
- The ST-2000XM CCD camera is malfunctioning (it only produces black images).
- Installed the heating and AC unit. The unit sits outside the obs. There are
two hole approx 2-1/2 inches in diameter. One is a feed the other is a return.
- Shimmed Southeast side of the of the observatory platform. The platform had
settled more than an inch. This caused rotation problems for the dome as the
battery for shutter had to rotate up-hill.
- The SBIG ST2000-XM is not functioning.
- Batteries showing low voltage.
- Received the HAC unit for the observatory.
- Plans to move the batteries outside the observatory. This wall free up space
- Weather-stripping was installed on the top of the lower shutter to fill the
gap between the upper and lower shutters.
- The weather-cam server was hung and required a restart.
- The battery issue has been solved. It appears the manually operated 12vdc
oscillating fan and a CCD camera were left on. The equipment ended up draining
the battery packs (this took about 3 weeks). The batteries have been recharged
and the charge regulator appears to be functioning OK.
- The two 100 amp hour marine batteries have gone dead. The battery packs are
currently being charged via a AC power charger. This will allow the interior
lights, cameras, fans and other 12 vdc items that are powered from the batteries
to function. The next step will be to check the water level of the battery
packs and test the solar charge regulator to be sure that it is functioning.
- Replaced the upper shutter. The wider flange of the new design allows the
shutters to be closed from any dome rotation position without the upper and
lower shutters smashing into each other. The solar vent will be installed at
a later date as will the lower shutter. Stops for the open and close limit
switches had to fabricated as the upper shutter didn't have them. Also it was
necessary to remove some fiberglass from the rear of the upper shutter to
prevent it from binding on the domes outer rails.
- The Weather Server is off-line for scheduled maintenance. That means no local
wind, rain, lightning, cloud, humidity or temp data. The Underground Weather &
the Citizen Weather Program data are also effected. The system should be back
up in the next 24 hrs (or sooner).
- The upper and lower shutters have developed a problem that is preventing
them from closing without manual intervention. A new set are on order and
should ship from the Sirius factory the week of 9/10/07. Until the new
shutters are installed the observatory will remain off-line.
- A terabyte of disk storage was added to the observatory's network. The
network drive will be used mostly as a backup device and to house the
observatory's star catalogs.
- The timer for the routine within the Observatory control program that parks
the telescope was changed. It is was increase from 60 seconds to 90 seconds.
At 60 seconds depending on where the the scope was pointing and what the slew
speed was set to, the Paramount ME would not reach the park position.
- Installed ESD dust covers on all open 9 pin serial ports.
- Used the iPhone to remotely control the observatory while on a business trip
from Sunnyvale, Ca.
The iPhone was used as the interface while connected to AT&T's EDGE network
in order to:
Power-up the observatory, telescope mount and cameras.
Configure, align and point the telescope.
Set-up and initiate the autoguider.
Focus the camera.
Begin the 15 minute exposure of the Bubble Nebula (NGC7635)
Process the image.
Shut-down and secure the cameras, telescope and observatory.
- Installed the additional counterweight on the Paramount. The Cal-K solar
scope was temporarily removed until the Dec axis of the mount can be better
- The data logger and the Weather Envoy for the weather station showed up
today after repair by Davis Instruments. They replaced the Data logger with
refurbished unit and they updated the firmware. It took a while to get the unit
setup and running again but it is back on-line.
- Updated the Observatory Control Panel GUI. The update adds access for the
Pyxis camera rotation and the Vellman data logger (used to monitor solar panel
and battery voltage). Two new buttons were added to allow the dome to be homed
or parked from the GUI. There was also some general clean-up to the GUI's
internal code in how command windows and processes are handled.
- Connected the AO-7 (Adaptive Optics) unit to the SBIG ST-2000XM CCD camera
which is attached to 10" SCT OTA. The camera has not been re-aligned.
- The replacement stepper driver chip ordered from Digikey arrived today. The
installation of the chip restored the LazyFocus to full functionality.
- A thunder storm took out the Wireless Weather Envoy. The WWE relays data
from the weather station to the weather PC (which uploads the data to this
web page). So there won't be any weather data until the WWE is repaired. I
believe the data logger module is fried. The unit has been returned to the
manufacturer. I estimate about 10 to 14 days for the repair.
- The LazyFocus emitted magic smoke this evening. Fortunately I was in the
observatory when it occurred. Upon inspection of the device I found the Stepper
Motor Driver chip was damaged. The chip (TEA3718) will have to be replaced. A
new one was ordered from Digikey.
- Move the GStar CCD video cam to the left side of SV102APO. This was done
because the camera cables interfered with Pyxis cables.
- Installed Pyxis camera rotator. It was originally acquired for use one the
ST-2000XM but the play in the gear drive and its inability to hold the ST2K in
position when the power to the unit is off relegated it for use on the ST-402ME.
- Recently in an attempt to improve the system the cloud sensor software was
moved to the observatory PC. An FTP server was installed on the Weather PC and
a remote serial port server was also installed on the Weather PC. The idea was
to allow the ACP program to access the cloud sensor. For that the happen the
software needs to be running on the observatory PC. This configuration turned
out to be to unreliable. The link between the serial port on the observatory
and the on on the weather PC would not auto-reconnect if the connection was
lost. This would cause the ACP scripts to fail because the loss of weather
data to be seen as "Sky Unsafe" message.
- A storm blew the cloud detector off the roof. It had be secured with ty-wraps.
Fortunately the detector landing on the rain soaked lawn so it wasn't damaged
in the fall. But it did get wet. It was dried, tested and re-secured using
stainless steel hose clamps.
- A fairly violent storm passed through area about 2:00 PM today. During the
storm the Cloud and Rain sensor was blown from the roof. The nylon ty-wraps
had became brittle due to Solar UV radiation. Fortunately the sensor landed
in the grass. It was dried out with a hair dryer become being remounted using
stainless steel hose clamps. It was also tested to be sure it still worked.
- Tweaked the Observatory Control Program (OCP8). It now full starts up and
shuts down the observatory. That includes homing both the dome and scope on
start-up. The shut-down procedure parks the dome, closes the shutters and
parks the scope.
- Last night a storm rolled through and had some lightning activity with it.
A strike took out com port 1 on the Weather PC and the wireless (DIY) shutter
control board. The control board was replaced with a spare. The serial port
on the Weather PC was connected to the wireless temperature and humidity sensor
located inside the observatory. The sensor still works and was moved to com2.
Com2 was connected to the wireless transmitter insider the house. Since a
transmitter was added to the observatory PC as part of the the ACP upgrade the
loss of the serial connection to inside transmitter won't result in any loss of
functions. The inside transmitter is still connected to the Boltwood sensor and
still functions the primary dome closure system in case of bad weather.
- The MaxDome II board has been installed and is operational.
ACP Dome Control update TO-Do-List
Purchase Maxdome-2 dome rotation controller.Done
This controller will replace the old Autodome controller currently used to
control the dome rotation. The new controller is ASCOM compatible.
Design wireless interface to Maxdome-2 controller board.Done
Needed for wireless control of dome rotation by special user programs.
Rewrite routines for JAT Observatory Control GUI.Done
Since the new wireless transmitter is located in the dome, the control program
will have to use a local serial port instead of the remote/virtual serial port for
the wireless transmitter. This includes the shutters, focuser switch box, and
the any custom software to manually move the dome.
Write stand alone control routines.Done
The stand alone routines are needed so ACP can control power and equipment.
These routines will have no user GUI. The following new routines are needed:
Top shutter open only (used when drift scanning).
Focuser switchbox control.
Dome rotation control (used by special user programs).
Install Clarity software for the Boltwood Cloud Sensor on the observatory PC.Done
Configure Network serial port.Done
The clarity software needs to access the Boltwood sensor which is not connected
to the observatory computer.
- Configure and test ACP control program In Process
Interface to MaxImDL 4.x
- Dark, Flats, Bias frames
> Scripts for light box control
- Interface to Pinpout
- Plate solving
- Mount pointing correction
Interface to Maxdome2
Remote user account setup
- The MaxDomeII Dome rotation controller arrived today. Now I just have to
install it and make it work.
- Clarity software for the Boltwood sensor is now running on the observatory
PC. That includes Russell Croman's Boltwood graphing program. I have not been
able to make the split serial port software work for the Clarity software so
the software on the Weather Station PC has been disabled. The Clarity software
in the observatory is connect to Com Port 6. Com6 is networked to com4 on the
- The ACP software is now using the Clarity software to determined the safety
of the sky. Anything below "Cloudy" is seen as an "Unsafe" Sky. Anything
above Wet or Very Cloudy is seen as a "Safe" sky. It would have been nice if
the ACP software would have displayed Clear, Cloudy, Very Cloudy of Wet like
the sensor software display does. It would have been more useful.
- Completed rewrite of remote control software GUI. The Observatory Control
Program is now at version 8. The software now uses the transmitter located in
the observatory and not the house. The house transmitter with still be used to
close the shutters when it rains. The Clarity software installed in the
observatory will eventually be linked to the ACP software. The stand alone
remote software OBS Start-up, Shut-down, Focuser#1 and Focuser#2 was also
rewritten to use the new transmitter.
- The OCP8 GUI now supports "Nudging the dome <<CCW and CW>>. The dome will
rotate for 1 second when the two new radio buttons are pressed.
- Created a place holder button on the GUI for the Adirondack Flat Field Light
- The Power Strip control was placed back onto Com2. The problem was the
CPS power strip needs to run @ 300 Baud, the serial port switch box will not
operate below 9600 so the 2 got not be used together.
- A Dynex USB to Serial adapter was installed in the observatory. It is
configured as serial port # 3.
- Installed LazyFocus control box. Made new "Y" Power cable
- Disconnected focusers from Paramount.
- Reconfigured Com2 for LazyFocus (Com2 was Power Strip)
- Configured Serial Switch to use slot "J" for Power Strip (now Com1)
- Installed wireless transmitter in observatory (Slot D on Serial Switch)
- Configured Serial Switch to use slot "G" for Flat Fielder Light Panel (Com1)
- Began writing software to control OBS side transmitter and account for new
- Reset Clear/Cloudy temp threshold on Boltwood Cloud sensor to -32 degrees F.
- Install solar powered red lamp on outside observatory near the door knob.
- All Astronomy/observatory software has been placed in it own case.
- Began storing install and setup software from weather & OBS PC onto CD-Rom
- Ordered Lazy Focus to allow the Meade micro-focusers to be controlled by
FocusMax while connected to the Paramount. Currently FocusMax can not talk
- working to figure out how to launch "exe" file from the ACP console and
browser. This is a requirement as this how observatory remote power-ups and
shut-downs will be done.
- Evaluating DC3 Dreams Scheduler 2.0
- Installed DC3 Dream's ACP 5.0 Observatory control software. The ACP software
required the 3.x version of MaxImDL be upgraded to 4.x. The problem is ACP will
not control the version of Sirius Dome Controller current in use by the
observatory. This will require an upgrade of the old Autodome hardware and
software. The remote browser has been assigned to port 30080.
- The disk provide by DC3 Dream is defective. I have requested a replacement.
- Collimated the SV102A with the ST-200XM and the F50 Finder with the
Mallincam to the 10" SCT with the ST-402ME.
- Found a number of screws loose on the dovetail plate and one dovetail adapter
thumb screw was loose for the SV102a.
- Noticeable hot pixels present on the GSTAR-EX CCD video camera with the
ambient temp being around 67 Degrees F.
- Modified the Observatory Control GUI to include control Mallincam and the
GSTAR-EX as well as the 20 port serial switch.
- The 20 port serial switch was installed in the observatory. The cable for
the video camera was installed and routed under the floor.
- All the video cables were covered in split-loom.
- The problem with the serial port switch not responding when the RS232 to
RS485 converter was connected has been solved. The problem was twofold. The
first issue is pins 2 & 3 on the RS232 side needed to be reversed when connected
to the switch so a null modem was install on each of the 2 converters. The
second issue is the RS485 side of the converter needs external power when
connected to the switch. Once these 2 conditions were satisfied the switch
worked as adverized.
- A cable was fabricated using 6 wire telephone line, although only 2 wires are
need on the camera side for control (the +485 and -485). 2 of the wires were
used for the Mallincam and 2 were used for the GSTAR-EX settings control. I
found the ground wires were not needed.
- Building of the hyper and thermal electric cooler remote wireless controller
for the Mallincam has been put on hold temporarily.
- Fabricated a cable to connect the GSTAR-EX video to the RS232 to RS485
GSTAR-EX camera.. The cable with be used to remote control the
- A 20 port serial switch was added to the observatory. This was done because
of the increasing serial port requirements. While the switch adds 20 more ports
only one port can be use at a time. The switch will be used to switch between
the flat field light box, wireless controller, video cameras controllers and
other RS232 devices. Currently the switch is hanging-up when the RS-232 to RS485
converters are attached to any of the port. The manufacturer of the switch has
- Removed the Mallincam color video camera from the observatory in order to
build an remote controller to operate the 7 and 14 second Hyper switches and
the thermal electric cooler switch. The controller was designed using relays
that will be connected to a wireless receiver. A new Beta GUI for the
Observatory Control program is already being coded. The camera is being sent
back to the vendor in Canada for an upgrade to the hyper system. The new system
allows 7, 14, 28 & 56 second frame integration. The upgrade add a new switch the
camera so the remote interface and software design won't be completed until
the returns from the vendor. The remote modification to the camera will
requires a hole be drilled in the camera's case and small cable and connector
will the attached. The plan is to still keep the functionality of the manual
switch while adding the remote control ability. The new Beta GUI will also
- Updated the observatory PC's mememory to 1 gig.
capture unit was installed in order to capture the output of the Mallincam. An
RCA composite video plug was connected (the S-video connector still goes to
- Tested the dome rotation. No problems and no noises. I suspect it was ice
in the track.
- Installed the GStar-Com remote control software on the observatory PC.
- Installed the Dazzle software. Dazzle is a video capture device that will
hardware will be installed tomorrow.
- Replacement 4 port USB hub will be installed in the observatory. The SBIG
cameras have been suffering from errors and non-responses.
- The dome started making a grinding noise when it was positioned southeast.
tested with the laptop. The interface is RS-485 and the GStar-EX software only
works with COM1. The observatory PC with be setup for attachment of the
MallinCam on COM1.
- Updated the Observatory and Weather station PC to Service Pack 2
to be powered from a single power supply.
- Installed Meade LPI
- F6.3 focal reducer removed from 10" SCT. The FL is increased from 1600mm to
2500mm @ F10.
- Shutter battery has been recharged.
- The voltage in battery for the shutters dropped the ~5 volts. This prevented
the top shutter from being closed. The emergency external supply was connected
to the battery in order to close the shutter. The observatory has been taken
off-line and the trickle charger was connected. I believe the reason for the
dead battery is extended cloudy weather. When the sky is cloudy the cloud
sensor continuously sends the closure command to the shutter receiver. Since
the battery is charged by a solar panel, if it remains cloudy for many days in
a row the battery will be drained.
- The 4.5 vdc external power connector for the wireless weather module in the
observatory managed to work itself loose. The module switch over to it's
internal batteries (which went or where dead) and stopped working. New
batteries where installed and the module was reconnected. The time and date was
reset. About 24 hours or weather data was lost.
- Installed a GStar-Ex B&W video camera with a Tamaron wide-angle on the side
of the SV102A. The camera is attached to an inverted Losmandy camera mount.
The camera mount is mount the dovetail plate on the top of the SV102A. The
camera is actually mount further forward than shown in the image. This was done
so the mounting clamps of the SV102a with be in the camera's view. The camera
is used to monitor the dome's slit in relation to the telescopes' aperture.
The output from the GStar-Ex is routed to the Slingbox A/V the same as the
Mallincam color video camera. The Mallincam is connected via a S-Video cable
and the GStar-Ex is connected via the RCA jack. This setup allows the video
output from from either camera to be selected remotely using the SlingPlayer
- Given up on controlling the Mallincam remotely (at lest for the time being).
The Mallincam has been mounted to the modified Stellarvue F50 finder scope.
The F50 has had its diagonal removed and a threaded 2" extension barrel has
been bonded to the removable back plate on the F50. The new configuration will
be used as an electronic video finder scope.
- Built 2 different cable in an attempt to control the Mallincam settings
remotely. Neither cable works. One is based on a RS-485 protocal the other is
based on RS-232. The software use is from a GStar-EX camera which it appears
the Mallincam is based on.
- Installed the Mallincam video camera. The camera is connected to the a
Slingbox AV unit. This allows the video to be accessed from any computer that
has the SlingPlayer software installed.
- Replaced the screw for the screw-drive for the lower shutter. The dome is
- The dome has been taken off-line to prevent remote operation. The reason is
a screw was found on the floor of the observatory. After a lengthy inspection
of the telescopes, pier and dome the location of where the screw came from
was found. The screw holds the screw-drive for the lower shutter and allows it
the pivot during opening and closing. Before the screw is replaced
non-permanent locking compound will be applied in order to prevent the screw
from working loose in the future.
- Reconfigured the SV102APO to use the stock Stellarvue mounting rings instead
of the 3 point Losmandy rings. The stock rings hold the scope more rigidly then
the 3 point rings. This should help eliminate most of the remaining flexing,
but will require shimming in order to align the pointing of the piggy-backed
scope with the main OTA. It will also no longer be possible to adjust the
piggy-backed scope in order to locate a guidestar if needed.
- Installed a CPS Intelligent Auto Push Board and a N-RRC controller to the
Weather Station PC. The PC is now able to be shut-down and booted remotely and
reliably. (This setup is identical to the Observatory PC setup).
- Added a link on website's Observatory status page for remote power control of
both the Observatory and Weather Station PC.
- Installed new fully charged shutter motor battery.
*** After 5 days FTP access to the website has finally been restored. The FTP
server for the observatory site stopped working @ 5:15 am Friday 11/10/06.
That meant no weather, lightening, cloud, sky-cam or weather-cam pictures
could be uploaded to the website, nor could the website be updated via FTP
during that time.
There is still no FTP access to the root directory and the FTP account name
in use before the outage is no longer valid, so all the scripts and automated
software that uploaded data to the website had to be changed.
So if you happen to be looking for web hosting services I would have to
recommend that you steer clear of IPOWER Web. In my opinion they suck.
- Created new local horizon mask. The mask defines what areas of the sky are
visible to the telescopes.
- I decided my ISP (IpowerWeb) sucks. There has been no FTP access to site
since Friday morning at about 5:15am. The CSR claims there was a power failure.
and the JATObservery domain was on one the FTP servers affected by the power
outage. Because of this issue no weather data, lightning data , cloud data or
images can be uploaded to the website.
- Ordered a new shutter battery. The new battery has slightly more capacity
than the original Panasonic battery supplied (9ah vs the original 7.5ah)
TR9-12 - TEMPEST SEALED LEAD-ACID AGM BATTERY
- Received the Baader solar filter for the Stellarvue SV102APO.
- The shutter battery is showing signs of not holding a good charge since the
temperature has begun to drop. The battery is currently being charged at night
by an AC charger. Because the charger is connected, the dome rotation has been
disabled so as not to allow remote start-up of the dome. Doing so with the
charging cable is connected will cause damage.
- Modified a Meade 8x50 finder scope to accept standard eyepieces.
- Aligned scope using the iterative method
- Installed an RJ-11 "Y" on the end of the 9 pin to RJ-11 connector on the
- The 12 volt trickle charger has been connected the the shutter battery.
- Installed a CPS Intelligent Auto Push Board into the Observatory PC. The PC
is now able to be shut-down and booted remotely and reliably. The Relay box
has been removed.
- Removed Windows Service Pack 2 update from the Observatory PC. The SBIG
ST-2000XM USB camera would not connect. After the SP2 was removed the camera
is functioning fine.
- Re-installed side cover on Paramount. Scope only eyeballed aligned to Dec 40
degree tick mark. The scope is operational but pointing will not be accurate.
- Remove the Windows Service Pack 2 update from the Weather Station PC. It was
hanging the WeatherLink software and preventing it from uploading correctly.
After the removal of Windows SP2 the software is functioning correctly.
- The free version of "LogMeIn